# Dynamics of the unitary Bose gas near a narrow Feshbach resonance: universal coherent atom-molecule oscillations

### v1

@@ Abstract
Introduction. Understanding the inherently unstable unitary Bose gas has remained a challenge  Papp et al. (2008); Navon et al. (2011); Pollack et al. (2009); Piatecki and Krauth (2014); Wild et al. (2012); Braaten et al. (2011). Some progress, however, has been made principally because rapid field sweeps across a Feshbach resonance to unitarity show, fortuitously, that the gas lives long enough to reveal features of quasi-steady state behavior before inevitable losses set in. This is seen through the momentum distribution $n(k)$ of non-condensed particles  Makotyn et al. (2014); Eigen et al. (2018). Importantly, these sweeps establish that the signal at high momentum $k$ grows and eventually saturates, as a function of time. These intriguing saturation phenomena have been referred to as a form of prethermalization Yin and Radzihovsky (2016); Rançon and Levin (2014). Interestingly, for the systems studied thus far, the time scale for such quasi-steady state dynamics is claimed to be universal. As the scattering length $a_{s}\rightarrow\infty$, it is presumed that time dependent phenomena should be determined by a single energy scale $E_{F}\propto(6\pi^{2}n)^{2/3}$ Cowell et al. (2002); Song and Zhou (2009); Lee and Lee (2010); Diederix et al. (2011); Li and Ho (2012); Jiang et al. (2014), where $n$ is the density of bosons.
@@
Figure 1: Effective temperature of atoms (black solid line) and molecules (blue) based on an error-bar weighted fit to data from Ref.Zhang et al. (2023) emphasizing the creation of out-of-condensate particles deriving from an atomic BEC after a quench to unitarity, see Supplement sup . The small cartoons indicate the different species of particles generated at different stages of the dynamics, including atomic and molecular BECs, finite-momentum atomic and molecular pairs.
In this paper, using both theory and experiment, we investigate the case of this narrow Feshbach resonance. In contrast to previous Bose gas literature which focused on the atoms, one has access to a sizeable population of closed channel molecules which govern much of the physics at unitarity. Our work is motivated by an earlier observation of the effective temperature, $T_{\rm{eff}}$ of out-of-condensate atoms and of closed channel molecules Zhang et al. (2023) which is plotted in Fig. 1. These are created during a quench to unitarity of an atomic Bose-Einstein condensate (BEC) of 133Cs atoms associated with a narrow resonance (of width 8.3 mG). Although previously oscillations in $T_{\rm{eff}}$ were not reported, they are revealed here based on an error-bar weighted fit to the temperature data.
It is our goal to experimentally investigate and theoretically understand the initial creation and subsequent oscillatory dynamics of these excited atomic and molecular states. To this end, we employ a time-of-flight (TOF) procedure. By analyzing a data set from previous experiments Zhang et al. (2023); Wang et al. (2024), we are able to characterize multiple new observables. These include the atomic and molecular particle number distribution at different momenta and their associated kinetic energies sup .
Theoretically to understand the oscillatory behavior observed in these quantities, we go beyond the Hartree-Fock-Bogolibov (HFB)  Kokkelmans and Holland (2002); Snyder et al. (2012) approach and establish a new dynamical framework that includes atom-molecule correlations. These correlations, which have no counterpart in the broad resonance case, turn out to be central to the dynamics of the out-of-condensate particles. In the process we identify three distinct evolutionary stages in the quench dynamics; this is initiated by the generation of a molecular BEC and subsequently followed by the creation of finite-momentum atoms and molecules which then undergo coherent oscillationsZhang et al. (2023); Tian et al. (2024). Satisfactory quantitative agreement is obtained between theory and experiment both for the rather large oscillatory frequency which turns out to be around $2$ kHz and for the net increase of kinetic energy (associated with the decay of the initial atomic condensate) which is transferred to newly excited atoms and molecules.
@@
Here $\sigma=1,2$ represents the atoms and molecules, respectively, and $\nu_{\sigma}=\delta_{\sigma,2}\nu$ where $\nu$ is the detuning; $g_{\sigma}$ is proportional to the background $s$-wave interaction, and $\alpha$ corresponds to the important Feshbach coupling strength between the two channels. Model parameters are related to the physical quantities by $g_{\sigma}=4\pi\hbar^{2}a_{\sigma}/(\Gamma_{\sigma}m_{\sigma})$, $\alpha=\Gamma_{1}^{-1}\sqrt{2\pi\hbar^{2}a_{1}\delta\mu\Delta B/m_{1}}$ and $\Gamma_{\sigma}$ is $1-2a_{\sigma}\Lambda/\pi$ where $\Lambda$ is a regularization cut-off. Here $a_{1}$ and $a_{2}$ are the background scattering length of atom and molecules, $\delta\mu$ is the relative magnetic moment between open and closed channels and $\Delta B$ is the resonance width. The detuning $\nu$ is related to the physical detuning $\nu_{r}=\delta\mu\delta B$ by $\nu=\nu_{r}+\Lambda m_{1}\Gamma_{1}\alpha^{2}/(\pi^{2}\hbar^{2})$ where $\delta B$ is the distance of the magnetic field from the resonance. We use the experimental values Zhang et al. (2023) for the parameters $a_{1},a_{2},\delta\mu,\Delta B$ and the initial atomic density $n$ in the theoretical calculations sup .
To theoretically study non-equilibrium dynamics, we focus on one and two-point equal-time correlations as dynamical variables. This approach is related to others in the literature Colussi et al. (2020); Van Regemortel et al. (2018) but now generalized to include the complexity of the two channel system. The one-point correlation $\xi_{\sigma}(t)=V^{-1/2}\langle\hat{\psi}_{\sigma}(k=0,t)\rangle$ represents the atomic or molecular condensate wavefunction, and $c_{\sigma}\equiv|\xi_{\sigma}(t)|^{2}$ represents the condensate density. Operators representing the excitations are then given by $\hat{\psi}_{\sigma}^{\prime}(k)=\hat{\psi}_{\sigma}(k)-\delta_{k,0}\sqrt{V}\xi_{\sigma}(t)$.
@@
The dynamical equations associated with these correlation functions are readily derived as
@@
Here $n_{\sigma}=\sum_{k\neq 0}n_{\sigma}(k)/V$ is the density of out-of-condensate atoms and molecules and $V$ is the volume. We define $x_{1}=\sum_{k\neq 0}G_{11}(k)/V$ and $x_{2}=\sum_{k\neq 0}G_{33}(k)/V$ which represent atom and molecule pairs, respectively, as discussed in the supplement sup . Additionally $f_{12}=\sum_{k}G_{23}(k)/V$ characterizes the atom-molecule correlation function which plays an important role. The equations of motion for the 2-point correlations can be compactly written as
@@
where $L(k)$ represents the coupling between finite-momentum atoms and molecules sup . This dynamical scheme should be contrasted with the standard Hartree-Fock-Bogoliubov approach Kokkelmans and Holland (2002); Snyder et al. (2012) which omit these new atom-molecule correlations including $f_{12}$.
@@
Figure 3: Kinetic energy variation $E_{K}/N_{0}$ for atoms and molecules which is associated with the population transfers and growth in Figure 2. Here $E^{\sigma}_{K}/N_{0}\equiv\int d^{3}k/(2\pi)^{3}n_{\sigma}(k)(k_{x}^{2}+k_{y}^{2})/(2m_{\sigma}n)$. Theory (top panel) and experiment (bottom panel). In experiment (but not theory) there is loss of particles which is associated with a gain in kinetic energy, but the two effects tend to at least partially compensate in the figure. Experiment and theory show similar oscillatory behavior and both show a net increase of kinetic energy which is of the order of $5$ nK times $N_{0}$, where $N_{0}$ represents the total number of particles at the start of the sweep.
Three Evolutionary Stages Using these dynamical equations we identify different stages of particle creation throughout the quench dynamics. The quench of an atomic BEC to unitarity is implemented by an abrupt change in the detuning $\nu_{r}\sim 0$ assuming initially, at $t=0$, all particles reside in an atomic BEC. Fig. 2 indicates the three distinct dynamical stages where different species of particles are generated. In stage I, Feshbach couplings convert the initial atomic condensate into a coherent molecular BEC along with a small number of atomic and molecular pairs having momentum, ±$\pm k$. 111 The growth of this molecular BEC is associated with instanton dynamics: $\xi_{2}(t)/\sqrt{n}=i\tanh(t/t_{\alpha})$, where $t_{\alpha}=\hbar/\alpha\sqrt{2n}$ sup . In stage II, shortly after its formation, the molecular BEC begins to partially decay and more atomic pairs appear. These latter in turn combine with an atom from the condensate to form an atom-molecule complex. The combination of an atom-molecule complex and an atom from the condensate is further converted into a pair of molecules. In stage III, with these finite-momentum atomic and molecular populations now fully formed, the system enters a stage of steady-state coherent oscillations.
@@
Most of the general observations we outline below (unless indicated otherwise) pertain to both theory and experiment 222 Because the experimentally observed particle loss is not significant within the 3 $\mathrm{ms}$ time frame Zhang et al. (2023), it makes sense to consider these comparisons. and these can be seen in Fig. 3. As is suggested by Figure 1, we find the newly generated out-of-condensate particles carry kinetic energy that also undergoes oscillatory dynamics. As the atomic condensate decays, a portion of the interaction energy is transferred to non-condensed particles as kinetic energy. This leads to an increase of the kinetic energy in both atoms and molecules right after the quench at $t<0.5~{}\mathrm{ms}$. We can quantify this increase, $E^{\sigma}_{K}/N_{0}$, which is of the order of 10 $\mathrm{nK}$ for the atoms and 5 $\mathrm{nK}$ for the molecules, as found in both theory and experiment.
After this initial rise, the kinetic energy subsequently oscillates and these oscillations persist for a period of around 2 $\mathrm{ms}$ notably, without any sign of thermalization. This oscillatory frequency is around 2 kHz, which, importantly, coincides with that of the population oscillations (see Fig. 2Zhang et al. (2023).) Indeed, this single oscillation pattern observed in all collective observables reflects the fact that the same underlying microscopic dynamical processes are present.
Momentum-resolved Distributions and Asymptotic Coherent Oscillations. Of interest next is to clarify the microscopic mechanism responsible for the generation of finite-momentum particles, after the initial formation of the molecular BEC. This is associated with the Feshbach resonant interaction and is contained in the properties of the matrix $L(k)$ in Eq. 3. The eigenvalues of $L$ determine the speed of growth of these out-of-condensate particles while the eigenvectors determine the fraction of atoms and molecules involved in the generation process. We consider the eigenvalues of $L(k)$,
@@
All of this can be contrasted with the larger $k$ regime where the eigenvalues are real and there is consequently no significant occupation of higher momentum states, say $k$ of the order of $k_{\mathrm{F}}$. Here the kinetic oscillations rather than the exponential growth dominates. Indeed, the momentum distributions we observe are consistent with the above analysis, where particles are seen to be concentrated in states with low momentum up to $\approx k_{F}/2$ where $k_{F}\equiv(6\pi^{2}n)^{1/3}$. We stress that, in contrast to the saturation effects seen in the broad Feshbach resonanceMakotyn et al. (2014); Eigen et al. (2018), the momentum distributions in the narrow resonance case oscillate during the later time dynamics, as seen in Fig. 4a,c.
@@
Conclusion. We have emphasized throughout that the coherent oscillations we observe for the narrow resonance case are universal as they are only determined by a single energy scale $\alpha\sqrt{n}$. This leads to a different paradigm for the universal dynamics of the unitary Bose gas. This is associated with a many body (Feshbach) coupling and is to be contrasted with the universality previously discussed in the literature  Makotyn et al. (2014); Eigen et al. (2018) and associated with the Fermi energy. As in this early literature, arriving at this universality requires us to focus on characterizing the dynamical behavior of non-condensed particles which appear when an atomic condensate is swept to unitarity. Here in addition to atoms there are out-of-condensate molecules, and once formed, all populations reach a steady oscillatory state; all finite-momentum particles are seen to have the same intrinsic oscillation frequency which is the same as that of the condensate Wang et al. (2024).
@@
Acknowledgement This work is supported by the National Science Foundation under Grant No. PHY1511696 and PHY-2103542, and by the Air Force Office of Scientific Research under award number FA9550-21-1-0447. We thank Cheng Chin for helpful discussions. Z.Z. acknowledges the Bloch Postdoctoral Fellowship. S.N. acknowledges support from the Takenaka Scholarship Foundation. Z.W. is supported by the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0301904). We also acknowledge the University of Chicago’s Research Computing Center for their support of this work.

### v2

# Universal coherent atom-molecule oscillations in the dynamics of the unitary Bose gas near a narrow Feshbach resonance

@@ Abstract
Introduction. Understanding the inherently unstable unitary Bose gas has remained a challenge  Papp et al. (2008); Navon et al. (2011); Pollack et al. (2009); Piatecki and Krauth (2014); Wild et al. (2012); Braaten et al. (2011). Some progress, however, has been made principally because rapid field sweeps across a Feshbach resonance to unitarity show, fortuitously, that the gas lives long enough to reveal features of quasi-steady state behavior before inevitable losses set in. This is seen through the momentum distribution $n(k)$ of non-condensed particles  Makotyn et al. (2014); Eigen et al. (2018), which are associated with nonzero momentum. Importantly, these sweeps establish that the signal at high momentum $k$ grows and eventually saturates, as a function of time. These intriguing saturation phenomena have been referred to as a form of prethermalization Yin and Radzihovsky (2016); Rançon and Levin (2014). Interestingly, for the systems studied thus far, the time scale for such quasi-steady state dynamics is claimed to be universal. As the scattering length $a_{s}\rightarrow\infty$, it is presumed that time dependent phenomena should be determined by a single energy scale $E_{F}\propto(6\pi^{2}n)^{2/3}$ Cowell et al. (2002); Song and Zhou (2009); Lee and Lee (2010); Diederix et al. (2011); Li and Ho (2012); Jiang et al. (2014); Gao et al. (2020), where $n$ is the density of bosons.
@@
Figure 1: Effective temperature of atoms (black solid line) and molecules (blue) based on an error-bar weighted fit to data from Ref.Zhang et al. (2023) emphasizing the creation of out-of-condensate particles deriving from an atomic BEC after a quench to unitarity, see Supplement sup . The small cartoons indicate the different species of particles generated at different stages of the dynamics, including atomic and molecular BECs, and finite-momentum atomic and molecular pairs.
In this paper, using both theory and experiment, we investigate the case of this narrow Feshbach resonance. In contrast to previous Bose gas literature which focused on the atoms, one has access to a sizeable population of closed channel molecules which govern much of the physics at unitarity. Our work is motivated by an earlier observation of the effective temperature, $T_{\rm{eff}}$ of out-of-condensate atoms and of closed channel molecules Zhang et al. (2023) which is plotted in Fig. 1. These are created during a quench to unitarity of an atomic Bose-Einstein condensate (BEC) of 133Cs atoms associated with a narrow resonance (of width 8.3 mG). Although previously oscillations in $T_{\rm{eff}}$ were not reported, they are revealed here based on an error-bar weighted fit to the temperature dependent data.
It is our goal to experimentally investigate and theoretically understand the initial creation and subsequent oscillatory dynamics of these excited atomic and molecular states. To this end, we employ a time-of-flight (TOF) procedure. By analyzing a data set from previous experiments Zhang et al. (2023); Wang et al. (2024a), we are able to characterize multiple new observables. These include the atomic and molecular particle number distribution at different momenta and their associated kinetic energies sup .
Theoretically to understand the oscillatory behavior observed in these quantities, we go beyond the Hartree-Fock-Bogolibov (HFB)  Kokkelmans and Holland (2002); Snyder et al. (2012) approach and establish a new dynamical framework that includes atom-molecule correlations. These correlations, which have no counterpart in the broad resonance case, turn out to be central to the dynamics of the out-of-condensate particles. In the process we identify three distinct evolutionary stages in the quench dynamics; these are initiated by the generation of a molecular BEC and subsequently followed by the creation of finite-momentum atoms and molecules which then undergo coherent oscillationsZhang et al. (2023); Tian et al. (2024); Sadhasivam et al. (2024). Satisfactory quantitative agreement is obtained between theory and experiment both for the rather large oscillatory frequency which turns out to be around $2$ kHz and for the net increase of kinetic energy (associated with the decay of the initial atomic condensate) which is transferred to newly excited atoms and molecules.
@@
Here $\sigma=1,2$ represents the atoms and molecules, respectively, and $\nu_{\sigma}=\delta_{\sigma,2}\nu$ where $\nu$ is the detuning; $g_{\sigma}$ is proportional to the background $s$-wave interaction, and $\alpha$ corresponds to the important Feshbach coupling Duine and Stoof (2004) strength between the two channels. Model parameters are related to the physical quantities by $g_{\sigma}=4\pi\hbar^{2}a_{\sigma}/(\Gamma_{\sigma}m_{\sigma})$, $\alpha=\Gamma_{1}^{-1}\sqrt{2\pi\hbar^{2}a_{1}\delta\mu\Delta B/m_{1}}$ and $\Gamma_{\sigma}$ is $1-2a_{\sigma}\Lambda/\pi$. Here $a_{1}$ and $a_{2}$ are the background scattering length of atom and molecules, $\delta\mu$ is the relative magnetic moment between open and closed channels and $\Delta B$ is the resonance width. The detuning $\nu$ is related to the physical detuning $\nu_{r}=\delta\mu\delta B$ by $\nu=\nu_{r}+\Lambda m_{1}\Gamma_{1}\alpha^{2}/(\pi^{2}\hbar^{2})$ where $\delta B$ is the distance of the magnetic field from the resonance. Here $\Lambda$ is a regularization cutoff, chosen here as $a_{1}\Lambda=\pi/10$, although the results are not sensitive to the specific numerical value. We use experimental values Zhang et al. (2023) for the parameters $a_{1},a_{2},\delta\mu,\Delta B$ to determine $g_{1},g_{2}$ and $\alpha$, and the initial atomic density $n$ in simulations sup .
To theoretically study non-equilibrium dynamics, we focus on one and two-point equal-time correlations as dynamical variables Colussi et al. (2020); Van Regemortel et al. (2018); van de Kraats et al. (2024). The one-point correlation $\xi_{\sigma}(t)=V^{-1/2}\langle\hat{\psi}_{\sigma}(k=0,t)\rangle$ represents the atomic or molecular condensate wavefunction, and $c_{\sigma}\equiv|\xi_{\sigma}(t)|^{2}$ represents the condensate density. Operators representing the excitations are then given by $\hat{\psi}_{\sigma}^{\prime}(k)=\hat{\psi}_{\sigma}(k)-\delta_{k,0}\sqrt{V}\xi_{\sigma}(t)$.
@@
The dynamical equations associated with these correlation functions are readily derived in terms of coupling constants $g_{1}$, $g_{2}$ and $\alpha$ which are, in turn, defined in terms of physical scattering lengths and the resonance width associated with the Hamiltonian of Eq. 1. These equations are given by
@@
Here $n_{\sigma}=\sum_{k\neq 0}n_{\sigma}(k)/V$ is the density of out-of-condensate atoms/molecules, which includes all finite-$k$ population of atoms/molecules, and $V$ is the volume. We define $x_{1}=\sum_{k\neq 0}G_{11}(k)/V$ and $x_{2}=\sum_{k\neq 0}G_{33}(k)/V$ which represent atom and molecule pairs, respectively, as discussed in the supplement sup . Additionally $f_{12}=\sum_{k}G_{23}(k)/V$ characterizes the atom-molecule correlation function which plays a centrally important role. The equations of motion for the 2-point correlations can be compactly written as
@@
where $L(k)$ represents the coupling between finite-momentum atoms and molecules sup . This dynamical scheme should be contrasted with the standard Hartree-Fock-Bogoliubov approach Kokkelmans and Holland (2002); Snyder et al. (2012) which omits these new atom-molecule correlations such as $f_{12}$.
@@
Figure 3: Kinetic energy variation $E_{K}/N_{0}$ for atoms and molecules which is associated with the population transfers and growth in Figure 2. Here $E^{\sigma}_{K}/N_{0}\equiv\int d^{3}k/(2\pi)^{3}n_{\sigma}(k)(k_{x}^{2}+k_{y}^{2})/(2m_{\sigma}n)$. Theory (top panel) and experiment (bottom panel). In experiment (but not theory) there is loss of particles which is associated with a gain in kinetic energy, but the two effects tend to at least partially compensate in the figure. Experiment and theory show similar oscillatory behavior and both show a net increase of kinetic energy which is of the order of $5$ nK times $N_{0}$, where $N_{0}$ represents the total number of particles at the start of the sweep. We focus here are the temperature variations which are comparable, noting that there are differences in the absolute values presumably because the experiments are at finite $T$, while the theory is at $T=0$. Thus, while panels (a) and (c) show some quantitative differences, it is reasonable that that the occupation of more and higher momentum states seen in experiment may reflect thermal effects; the experiments are performed at finite $T$, whereas theory is at $T=0$.
Three Evolutionary Stages Using these dynamical equations we identify different stages of particle creation throughout the quench dynamics. The quench of an atomic BEC to unitarity is implemented by an abrupt change in the detuning $\nu_{r}\sim 0$ assuming initially, at $t=0$, all particles reside in an atomic BEC. As indicated in Fig. 2 in stage I, Feshbach couplings convert the initial atomic condensate into a coherent molecular BEC along with a small number of atomic and molecular pairs having momentum, ±$\pm k$. 111 The growth of this molecular BEC is associated with instanton dynamics: $\xi_{2}(t)/\sqrt{n}=i\tanh(t/t_{\alpha})$, where $t_{\alpha}=\hbar/\alpha\sqrt{2n}$ sup . In stage II, shortly after its formation, the molecular BEC begins to partially decay and more atomic pairs appear. These latter in turn combine with an atom from the condensate to form an atom-molecule complex. The combination of an atom-molecule complex and an atom from the condensate is further converted into a pair of molecules. In stage III, with these finite-momentum atomic and molecular populations now fully formed, the system enters a stage of steady-state coherent oscillations.
@@
Most of the general observations we outline below (unless indicated otherwise) pertain to both theory and experiment 222 Because the experimentally observed particle loss is not significant within the 3 $\mathrm{ms}$ time frame Zhang et al. (2023), it makes sense to consider these comparisons. and a comparison is presented in Fig. 3. As is suggested by Figure 1, we see here that the newly generated out-of-condensate particles which carry kinetic energy are associated with oscillatory dynamics. As the atomic condensate decays, a portion of its interaction energy is transferred to non-condensed particles. This leads to an increase of the kinetic energy in both atoms and molecules immediately after the quench at $t<0.5~{}\mathrm{ms}$. We can quantify this increase, $E^{\sigma}_{K}/N_{0}$, which is of the order of 10 $\mathrm{nK}$ for the atoms and 5 $\mathrm{nK}$ for the molecules, as found in both theory and experiment.
After the initial growth period, the oscillations persist for a period of around 2 $\mathrm{ms}$ notably, without any sign of thermalization. This oscillatory frequency is around 2 kHz, which, importantly, coincides with that of the population oscillations (see Fig. 2Zhang et al. (2023).) Indeed, this single oscillation pattern observed in all collective observables reflects the fact that the same underlying microscopic dynamical processes are present.
Momentum-resolved Distributions and Asymptotic Coherent Oscillations. Of interest next is to clarify the microscopic mechanism responsible for the generation of finite-momentum particles, after the initial formation of the molecular BEC. This is associated with the Feshbach resonant interaction and is contained in the properties of the matrix $L(k)$ in Eq. 3. The eigenvalues of $L$ determine the speed of growth of these out-of-condensate particles while the eigenvectors determine the fraction of atoms and molecules involved in the generation processWang et al. (2024b). We consider the eigenvalues of $L(k)$,
@@
All of this can be contrasted with the larger $k$ regime where the eigenvalues are real and there is consequently no significant occupation of those higher momentum states, having $k$ of the order of, say, $k_{\mathrm{F}}$. Here the kinetic oscillations rather than the exponential growth dominates. Indeed, the momentum distributions we observe are consistent with the above analysis, where particles are seen to be concentrated in states with low momentum up to $\approx k_{F}/2$ where $k_{F}\equiv(6\pi^{2}n)^{1/3}$. We stress that, in contrast to the saturation effects seen in the broad Feshbach resonanceMakotyn et al. (2014); Eigen et al. (2018), the momentum distributions in the narrow resonance case oscillate during the later time dynamics, as seen in Fig. 4a,c.
@@
Conclusion. We have emphasized throughout that the coherent oscillations we observe for the narrow resonance case are universal as they are only determined by a single energy scale $\alpha\sqrt{n}$. This leads to a different paradigm for the universal dynamics of the unitary Bose gas. This is associated with a many body (Feshbach) coupling and is to be contrasted with the universality previously discussed in the literature  Makotyn et al. (2014); Eigen et al. (2018) and associated with the Fermi energy. As in this early literature, arriving at this universality requires us to focus on characterizing the dynamical behavior of non-condensed particles which appear when an atomic condensate is swept to unitarity. Here in addition to atoms there are out-of-condensate molecules, and once formed, all populations reach a steady oscillatory state; all finite-momentum particles are seen to have the same intrinsic oscillation frequency which is the same as that of the condensate Wang et al. (2024a).
@@
Acknowledgement This work is supported by the National Science Foundation under Grant No. PHY1511696 and PHY-2103542, and by the Air Force Office of Scientific Research under award number FA9550-21-1-0447. We thank Cheng Chin for very helpful discussions. Z.Z. acknowledges the Bloch Postdoctoral Fellowship. S.N. acknowledges support from the Takenaka Scholarship Foundation. Z.W. is supported by the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0301904). We also acknowledge the University of Chicago’s Research Computing Center for their support of this work.